The following linear regression model can be used to predict ticket sales at a popular water park (the correlation is significant).

Ticket sales per hour = -631.25 + 11.25(current temperature in °F)

What is the predicted number of tickets sold per hour if the temperature is 86°F?
Answer #1

Given that, the regression model is,

Tickets sales per hour= -631.25+11.25(current temperature in °F)

We want to predict the number of tickets sold per hour if the temperature is 86 °F.

=> Tickets sales per hour = -631.25 + (11.25 * 86)

=> Tickets sales per hour = -631.25 + 967.5

=> Tickets sales per hour = 336.25

Answer : 336.25